OpenStudy (anonymous):

A triangle with side lengths of 5, 12, and 13 is a right triangle.

OpenStudy (anonymous):

A right triangle is: a^2 + b^2 = C^2 Does 5^2 + 12^2 = 13^2 ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Then it is not a right triangle :D